Operations Report – September 9th

MDRS Complete Operations Report 9-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: The seal on the Hab toilet appears to be leaking.

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 54.7

Minimum 23

Maximum 100

Note on solar: Kept rovers plugged in but did not use my AC this evening. Will keep monitoring.

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 300 gallons

Static to Loft Pump used – yes
Water in GreenHab – 120 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 200 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: no
Hours: 147.2
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: no
Hours: 135.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: yes
Hours: 141.3
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): none
Reason for use: n/a
Oil Added? Yes, to 1 and 2. 3 has very dirty old.
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling power washer to town. We have used 600 gallons of water so far
CrewCar used and why, where? No, the spare tire storage issue resolved

Luna used and why, where? No

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Nothing new to report.

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: Due to the failure of the town’s septic system, we are currently unable to pump the main system. It is ¾ full and I will just monitor it for now. We should be able to pump sometime in October.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Septic tank was emptied.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: We spent the entire day cleaning the exterior of the building to prep for painting. As clay is a charged particle, it clings to the exterior and is very hard to remove. This is why over time the building has taken on a pinkish-red hue. There was a steep learning curve, but we are finally getting it done. By nightfall we were about 80 % done, having used almost 600 gallons of water. Tomorrow morning we will finish the dome, then move on to cleaning the Hab while allowing the mud to dry around the ScienceDome before we begin painting. The only damage we caused was a small hole on the southwest side, which we will patch. The ScienceDome is white again. We also found some areas where the original paint job was only a single coat, so we have decided to do two coats of paint at this time.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Still too hot.

Questions, concerns, supplies needed and requests: Nothing at this time.

Copyright © The Mars Society. All rights reserved. | Main Site